OpenStudy (anonymous):

Can someone tell me if i did this right. -2y - 7 > 11 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

OpenStudy (anonymous):

Part 1: -2y - 7 > 11 to make the equation simple, you can change the > sign to an = sign. -2y - 7 = 11 now you will add seven to both sides +7 +7 -2y = 18 now you will divide -2 to both sides -- -- -2 -2 now you get X X = -9 but now you put your > sign back into the equation. X > -9 Part 2: To graph this you would have an open circle on -9 shading to the right.

OpenStudy (anonymous):

ok thank you i had a feely something was wrong

OpenStudy (anonymous):

Whenever you multiply an inequality by -1, the direction of the arrow needs to be flipped, so y < -9

OpenStudy (anonymous):

I agree with rowan! should be y < -9

OpenStudy (anonymous):

thus to graph, open circle on -9 and shade to left!

OpenStudy (anonymous):

-2y-7>11 -2y>18 -y>9 (remember to flip sign) y<-9

OpenStudy (anonymous):

ok thank you